Super Mario is standing on a 4.0 m tall mushroom. When he jumps off of it, what will his velocity be right as he is about to hit the ground?


1
Expert's answer
2021-01-27T17:29:13-0500
"KE=PE,""mgh=\dfrac{1}{2}mv^2,""v=\sqrt{2gh}=\sqrt{2\cdot9.8\ \dfrac{m}{s^2}\cdot4.0\ m}=8.85\ \dfrac{m}{s}."